On 8/15/2019, a 3-year forward contract, expiring 8/15/2022, on a non-dividend-paying stock was entered into when the stock price was $50 and the risk-free interest rate was 10.5% per annum with continuous compounding. 1 year later, on 8/15/2020, the stock price becomes $57. What is the "delivery" price of the forward contract entered into on 8/15/2019?
Answer #1

Forward Price (delivery Price) can be computed with following formula:

F= So * et

where,

F = Forward Price

S0 = Spot price of stock on Initiation

r = risk free rate

T = Maturity

Putting the value:

F = 50 * 0.10543

F = 50 * 1.370259310957

F= 68.5129655478498

F = $68.51

Please note, Forward Price (delivery price) remain constant through out the Forward contract life and does not change. On maturity, A investor with long forward position buy the stock at given forward price and a investor with short position sale the stock at given forward price. However, Value of Forward changes with time.

Thus, In given case, Forward price would be $68.51

Hope this will help, please do comment if you need any further explanation. Your feedback would be highly appreciated.